# Tree networks of real-world data: analysis of efficiency and spatiotemporal scales

### v3

@@ 1 Introduction
Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully satisfactory answer. In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [2]. The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics of fields in network theory [1, 3, 4, 5]. In living systems, natural selection is the way the information about the environment flows into the system’s genome which will then encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system. Thus, the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6]. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10, 11] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines. Here we challenge this hypothesis. In our work, a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
@@
We take as a case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [13]. The algorithm employs a tree-formation protocol where raw data are reorganized into hierarchical tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing network. At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space. Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$. For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out. For some intermediate value of $r$ a ragged tree structure with progressively smaller distances between the patterns is produced, while for $r\rightarrow 2$ a 1d-lattice is created. The algorithm produces exclusively trees, i.e. loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ is the number of links and nodes respectively. Using the metaphor introduced by Jacob for natural selection [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
@@ 2.1 Statistical physics of networks
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [20, 22]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location at time $\tau$ [4]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
@@
where $\{\mu_{i}(\tau)\}_{i=1}^{N}$ is the set of $\hat{\rho}(\tau)$ eigenvalues that are related to the set of Laplacian eigenvalues through $\mu_{i}(\tau)=e^{-\tau\lambda_{i}}/Z$. The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes. Thus the von Neumann entropy $S$ as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [22]. Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
The network entropy thus counts the (logarithmic) number of informationally connected components on the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat) $C=-dS/d\log\tau$ it is possible to identify the characteristic resolution scales of the network [22, 3]. For instance, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network. The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network. Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the full network revealing a single trivial scale. On the contrary, a constant specific heat reflects the scale-invariance of the network in a resolution region where information is lost at a constant rate. Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed in a sufficiently extended scale-interval [3]. The latter sentence makes for a definition for the scale-invariant networks [23]. The specific heat plateau height, $c_{0}$, is also related to the spectral dimension of the network $d_{s}=2c_{0}$ [3, 23].
@@ 2.2 Thermodynamics of networks
The statistical physics theory for networks introduced above allows us to view network phenomena from a new perspective and possibly to bridge the gap to network macroscopic observables. To this regard network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges results in diminishing the entropy of the network. Where lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. The entropic cost is given by:
@@
On the other hand, adding edges increases the information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [1, 4]. Less obvious is how, for a fixed number of edges, the degree distribution influences the flow of information. As the Laplacian governs the continuous-time edge-centric random walk, the presence of large hubs implies an enhancement of the process with faster propagation of the diffusing quantity [16, 24]. As the free energy is inversely related to the partition function, $Z=\sum_{i=1}^{N}e^{-\tau\lambda_{i}}$, the bounds of the former are given by the bounds of the latter. From $1\leq Z\leq N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$. Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
@@ 3 Methods
Figure 2: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work. In green an estimate is given computing the Lempel-Ziv complexity [29], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [30, 31, 32]. In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each. The left axis is referred to the quantities in green and orange, while the right axis is referred to the multilayer-perceptron accuracy. We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
@@ 4 Phase portrait
Figure 5: Multiple phase transitions in hierarchical tree networks. In (a) the specific heat $C$ across the range $r\in[0,2]$. At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat. The scale-invariant region shows a second phase transition from amorphous to regular 1d-lattice at $r\simeq 1.05$ indicated by the dotted white line. For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot c_{0}$, where $c_{0}$ is the plateau height in (c). Note that at the transition point $d_{s}=4/3$. In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.25,~{}0.35,~{}0.40,~{}0.70,~{}0.90,~{}1.05,2.00\}$. Dashed blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively. In (d), (e) the modularity and the normalized diameter, $\hat{d}=d/m$, together with their respective variances (dashed lines) identify the transition points. The modularity resolves the transition involving the mesoscale structure, while the diameter resolves the transition at the larger network scale. The legend indicates the two network sizes $N=500,~{}1000$. In (f) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-scales, i.e. the number of nodes included in the scale-invariant region. Note the discontinuity at the transition from star to amorphous. The dotted line shows the peak of the variance at the transition point. In (g) the entropy surface is depicted and the entropy at the first and second specific heat peaks are illustrated, see legend. In (h) the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate. In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right are depicted illustrating the two phase transitions at the two resolution scales. All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere.
The Portegys algorithm produces tree structures as those depicted in panel (b) of Figure 5 depending on the value of the control parameter $r$. Two structural phase transitions involving two different resolution scales can be identified (see panel (a)). The transition at the largest resolution scale can be described using as an order parameter a quantity that describes the large-scale network properties: $\hat{d}=d/m$, where $d$ is the diameter of the tree and $m$ is the number of edges equal to $m=N-1$. For hierarchical trees, another choice would be the normalized number of hierarchical levels $\hat{h}=h/(N-1)$ where $h$ is the number of hierarchical levels ranging from $h\in[0,N-1]$ (see illustration in panel (b)). Both quantities describe the "volume" expansion controlled by the parameter $r$. We chose the normalized diameter $\hat{d}$ and we identify this phase transition at $r\sim 1.05$ as indicated by the peak of the variance (dotted lines) in panel (e) for the two network sizes $N=500,~{}1000$ as indicated in the legend. The network structure transits from the amorphous regime to the regular 1d-lattice. Just below the transition point, a spectral dimension of $d_{s}=4/3=2c_{0}$ characteristic of random trees [34] is measured as indicated in (a) and shown in panels (c). Above the transition $d_{s}=1$, as it is expected for a 1d-lattice. Where $c_{0}$ is the specific heat plateau height.
@@
The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$. Here, the emergence of scale-invariant structural properties is identified by the appearance of a second peak in the specific heat and a flat plateau in between the two $C$-peaks, as shown in panels (a) and (c) of Figure 5. Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. The emergence of a plateau $C=c_{0}>0$ for $r>0.35$ reflects the emergence of progressively larger scale-invariant structures that for $r\rightarrow 2$ would include the full network in the thermodynamic limit $N\rightarrow\infty$. This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies occur either in the vicinity of this transition border or are included in the scale-invariant regime. This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network. In fact, we observe in panel (e) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition. Nevertheless, a phase transition indeed exists and an appropriate order parameter needs to be employed to identify it.
@@
This phase change occurs at smaller resolution scales, specifically, the mesoscales involved in the structural transition from the star graph to the scale-invariant regime (see the bifurcation in the specific heat and the related quantities in panels (a,c,g,h)). Thus, we can identify this transition using as an order parameter the modularity [35] of the tree which we show in panel (e). We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36]. We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor. This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
@@
In panel (f), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks counts the normalized (log)number of informationally connected nodes across the $\tau$-scales where the entropy is consumed at a constant rate, i.e. across the specific heat plateau. In simple terms, we count the fraction of nodes included in the scale-invariant region. We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
@@
As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see panel (h)). For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase. Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [3, 23].

### v4

@@ 1 Introduction
Why complex structures emerge in a real-world environment is a fascinating question that has not yet found a fully satisfactory answer. In a recent article [1], the emergence of complex structures is devised as the macroscopic thermodynamic result of the optimal ability to store the information about the environment in an accessible manner with the ability to transmit information across the structure [2]. The latter is defined in terms of microscopic quantities thanks to the fairly new developments in statistical physics in network theory [1, 3, 4, 5]. In living systems, natural selection is the way the information about the environment flows into the system’s genome which will then encode such information in the structural traits of the organisms which in turn reflect the functional capability of the system. Thus, the structural traits of a system could themselves enclose the complexity of the environmental conditions the system needs to adapt to, or in other words, to represent into its very own structure [6]. How these traits are optimized to maximize the fitness to the environment is an open question. The hypothesis of optimal functioning at the border of order and disorder suggests the critical point as the potential candidate to obtain functional advantages [7, 8, 9, 10, 11] and maximize the fitness to the outer world, or to obtain a better performance in a statistical generalization problem in the context of learning machines. Here we challenge this hypothesis. In our work, a system seeks the best-fitness configuration given some external constraints $r$ and a high-dimensional vector collection, the dataset $\mathcal{D}$.
@@
We take as a case study the algorithm introduced by Portegys [12] for accelerating nearest-neighbor (1-NN) pattern recognition tasks and subsequently extended to k-NN [13]. The algorithm employs a tree-formation protocol where raw data are reorganized into hierarchical tree structures using a recursive tree-formation method in which samples are randomly selected and inserted into the growing network. At each step, the new pattern is linked to either the current parent pattern or to one of the k-NN children of the parent pattern. The choice is based on the distance between the parent and the new pattern and the one evaluated between the child pattern and the new pattern. The parameter $r$ modulates the relative importance of these two distances creating a larger or smaller region of influence of the parent in the sample space. Figure 1 shows the three structure prototypes produced depending on the value of the control parameter $r$. For $r\sim 0$, every node is linked to the root node (see the central node, which is the first node selected in the process) and a star graph turns out. For some intermediate value of $r$ a ragged tree structure with progressively smaller distances between the patterns is produced, while for $r\rightarrow 2$ a 1d-lattice is created. The algorithm produces exclusively trees, i.e. loop-less networks, which are necessarily sparse $|E|=N-1$ where $|E|$ and $N$ are the number of links and nodes respectively. Using the metaphor introduced by Jacob for natural selection [14], the algorithm does act as a tinkerer that creates a structure using what is available at the time, collocating one node after the other across the tree formation process according to a parameterized distance-based rule.
@@ 2.1 Statistical physics of networks
The elements $\rho_{ij}$ represent the normalized amount of information transferred in a diffusion process between nodes $i$ and $j$ at time $\tau$. Remarkably $\rho_{ij}$ takes into account all possible topological pathways between the two nodes and assigns smaller weights to longer ones tantamount to path-integrals [20, 22]. The partition function $Z=\operatorname{Tr}[\hat{K}]=\sum^{N}_{i=1}e^{-\tau\lambda_{i}(\hat{L})}$ is a function of the eigenvalues, $\lambda_{i}$, of the Laplacian and is proportional to the average return probability of a continuous time edge-centric random walker to be back in its initial location after time $\tau$ [4]. Note that the object $\hat{\rho}$ is a positive semi-definite and Hermitian matrix whose trace sums to unity which has allowed its eigenvalues to be interpreted as probabilities.
@@
where $\{\mu_{i}(\tau)\}_{i=1}^{N}$ is the set of $\hat{\rho}(\tau)$ eigenvalues that are related to the set of Laplacian eigenvalues through $\mu_{i}(\tau)=e^{-\tau\lambda_{i}}/Z$. The entropy $S(\hat{\rho}(\tau))$ is a function of the normalized time $\tau$ and is bounded between $[\log C_{conn},~{}\log N]$ where $C_{conn}$ is the number of connected components and $N$ is the number of nodes. Thus the von Neumann entropy, $S$, as a function of $\tau$ reflects the entropic phase transition of information propagation (or loss) over the network [22]. Specifically for a connected network, at $\tau\rightarrow 0$, $S(\tau)=\log N$ reflects the segregated heterogeneous phase where information diffuses from the single nodes to the very local neighborhood; at $\tau\rightarrow\infty$, $S(\tau)=0$ the diffusion is governed by the smallest non-zero eigenvalue of the Laplacian, associated with the so-called Fiedler eigenvector, reflecting the integrated homogeneous phase where the information has propagated all over the network.
The network entropy thus counts the (logarithmic) number of informationally unconnected nodes in the heterogeneous network and by the inspection of the entropy $S$ and its generalized susceptibility (or specific heat), $C=-dS/d\log\tau$, it is possible to identify the characteristic resolution scales of the network [22, 3]. For instance, a single large peak of the specific heat $C$ identifies a single scale where there is a fast and unique information diffusion (or information lost) across the network. The latter is the typical situation of a star graph, depicted in Figure 1, or an Erdös-Rényi network. Before the peak, the informationally connected components are the single nodes, while after the peak the informationally connected components are the full network revealing a single trivial scale. On the contrary, a constant specific heat reflects the scale-invariance of the network in a resolution region where information is lost at a constant rate. Such a situation is typical of homogeneous $d$-dimensional lattices and of heterogeneous networks whose spectral density is power-law distributed in a sufficiently extended scale-interval [3]. The latter sentence makes for a definition for the scale-invariant networks [23]. The specific heat plateau height, $C_{0}$, is also related to the spectral dimension of the network $d_{s}=2C_{0}$ [3, 23].
@@ 2.2 Thermodynamics of networks
The statistical physics theory for networks introduced above allows us to view network phenomena from a new perspective and possibly to bridge the gap to network macroscopic observables. To this regard network thermodynamic quantities such as work, $W$, and heat, $Q$, have been defined in mathematical analogy to their thermodynamic counterparts [1]. Since a network of isolated nodes is the only network with maximum entropy $S_{iso}=\log N$ for $\tau>0$, progressively adding edges diminishes the network’s entropy. Where lower entropy reflects the diminished ability of the network to express diversity to a stochastic perturbation, or in other terms, the loss in the internal representation capabilities encoded in the network topological traits. The entropic cost is given by:
@@
On the other hand, adding edges increases the gain in information flow quantified by the (Helmholtz) free energy $F=-\log Z/\tau$ [1, 4]. Less obvious is how, for a fixed number of edges, the degree distribution influences the flow of information. As the Laplacian governs the continuous-time edge-centric random walk, the presence of large hubs implies an enhancement of the process with faster propagation of the diffusing quantity [16, 24]. As the free energy is inversely related to the partition function, $Z=\sum_{i=1}^{N}e^{-\tau\lambda_{i}}$, the bounds of the former are given by the bounds of the latter. From $1\leq Z\leq N$, where the lower bound is given for the connected network at large $\tau\gg 1$, while the upper bound is given by a network of isolated nodes $\forall\tau$, or for a connected one at small scales $\tau=0$. Thus, the $-\log N/\tau\leq F\leq 0$ and the gain in information flow is reflected by the variation of the free energy from the case of isolated nodes:
@@ 3 Methods
Figure 2: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work. In green an estimate is given computing the Lempel-Ziv complexity [29], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [30, 31, 32]. In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each. The left axis refers to the quantities in green and orange, while the right axis refers to the multilayer-perceptron accuracy. We observe that the CIFAR10 dataset is the most complex one, while the NIST dataset is the least complex among the four.
@@ 4 Phase portrait
Figure 5: Multiple phase transitions in hierarchical tree networks. In (a) the specific heat, $C$, across the range $r\in[0,2]$. At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat. The scale-invariant region shows a second phase transition from amorphous to regular 1d-lattice at $r\simeq 1.05$ indicated by the dotted white line. For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot C_{0}$, where $C_{0}$ is the plateau height in (c). Note that at $r\sim 0.8$ the spectral dimension is $d_{s}=4/3$ as expected for random trees with finite variance in the node degree. In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.35,~{}0.40,~{}0.70,~{}0.80,~{}0.90,2.00\}$. Vertical blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively. In (d), the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate. In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right. All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations.
Figure 6: Order parameters. In (a), the modularity, $Q$, resolves the transition involving the mesoscale structure. In (b), the normalized diameter, $\hat{d}=d/m$ resolves the large-scale transition. Their respective variances (dashed lines) identify the transition points. In (c), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-domain, i.e. the number of nodes included in the scale-invariant region. Note the discontinuity at the transition from star to amorphous. Curves in (a-b) are realized for networks of $N=500,~{}1000$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere. Curves in (c) are for networks of $N=500$.
[…]
The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$. Here, the emergence of scale-invariant structural properties is reflected by the progressively larger specific heat plateau in between the two $C$-peaks, as shown in (a) and (c) of Figure 5. For $r\rightarrow 2$, the plateau would include the full 1d-lattice in the thermodynamic limit $N\rightarrow\infty$.
Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. After the phase transition, a second peak of the specific heat emerges at larger $\tau$ scales. This second peak of $C$ in Figure 5.(c) identifies the scale at which the information is integrated between the distinct modules. In other words, the second peak reflects the "degree" of the large-scale modules, in much analogy to the first peak, which reflects the presence of hubs at the local connectivity scale of the nodes. Note that we exclude both the clusters, i.e. triangles, and higher-order cycles from our discussion as the tree networks we address are purely loop-less.
@@
This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies always occur in the scale-invariant region of the phase plane, and in some cases, they occur in the vicinity of this transition border $r\sim 0.35$. This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network. Indeed, we observe in Figure 6.(b) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition. Nevertheless, a phase transition exists and an appropriate order parameter needs to be employed to identify it.
@@
This phase change occurs at smaller resolution scales with respect to the first phase transition we discussed. Specifically, it involves the mesoscales properties changing from the star graph to the scale-invariant regime (see the bifurcation in the specific heat in Figure 5.(a),(c)). Thus, we can identify this transition using as an order parameter the modularity [35], $Q$, of the tree which we show in panel (a) of Figure 6. We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36]. We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor. This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
@@
As a second possible order parameter, for the mesoscale phase-transition, we show in Figure 6.(c) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks. This order parameter counts the normalized (log)number of informationally connected nodes across the $\tau$-scales of the specific heat plateau. In simple terms, we count the fraction of nodes included in the scale-invariant region. We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
@@
As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see Figure 5.(d)). For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase. Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [3, 23].